# New versions of Frobenius and integral closure of ideals

### v2

@@ 3.4. Comparing the new closures
Later in Section 4 (Proposition 4.12), we see that the inclusions above are generally strict. We note that each of the adjusted closures are insensitive to replacing the base ideal $I$ with its corresponding unadjusted closure.
@@
Let $I,\mathfrak{a}\subset R$ be ideals. Then, we have the following equalities.
@@
    $(I^{F})^{F\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$ and ($\,{}^{\mathfrak{a}^{t}}(I^{F})^{F}=\,^{\mathfrak{a}^{t}}I^{F}$, and
@@
    $(I^{*})^{*\mathfrak{a}^{t}}=I^{*\mathfrak{a}^{t}}$ and ($\,{}^{\mathfrak{a}^{t}}(I^{*})^{*}=\,^{\mathfrak{a}^{t}}I^{*}$, and
@@
    $\overline{J}^{\mathfrak{a}^{t}}=\overline{I}^{\mathfrak{a}^{t}}$, where $J=\overline{I}$.
@@
Since (${}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}={}^{\mathfrak{a}^{t}}I^{F}$ and (${}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{*})^{*}={}^{\mathfrak{a}^{t}}I^{*},$ it is easy to check the equalities using the inclusions $I\subset I^{F}\subset{}^{\mathfrak{a}^{t}}I^{F}$ and $I\subset I^{*}\subset{}^{\mathfrak{a}^{t}}I^{*},$ so we turn to the Hara-Yoshida versions.
@@
    We will show $(I^{F})^{F\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$. As $I\subset I^{F}$, we have $I^{F\mathfrak{a}^{t}}\subset(I^{F})^{F\mathfrak{a}^{t}}$. To see the reverse containment, let $x\in(I^{F})^{F\mathfrak{a}^{t}}$. Then $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{F})^{\left[p^{e}\right]}=I^{\left[p^{e}\right]}$ for all $e$ large, in particular, for $e$ at least the Frobenius test exponent [^1] There is an $e$ so that $(I^{F})^{\left[p^{e}\right]}\subset I^{\left[p^{e}\right]}$, the least such $e$ is called the Frobenius test exponent of $I$. See [Mad19] for survey of Frobenius test exponents. of $I$. Thus, $x\in I^{F\mathfrak{a}^{t}}$.
@@
    We will show $(I^{*})^{*\mathfrak{a}^{t}}=I^{*\mathfrak{a}^{t}}$. As $I\subset I^{*}$, we have $I^{*\mathfrak{a}^{t}}\subset(I^{*})^{*\mathfrak{a}^{t}}$. To see the reverse containment, let $x\in(I^{*})^{*\mathfrak{a}^{t}}$. Then there exists $c\in R^{\circ}$ such that $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{*})^{\left[p^{e}\right]}$ for all $e$ large. There is a $d\in R^{\circ}$ such that $d(I^{*})^{\left[p^{e}\right]}\subset I^{\left[p^{e}\right]}$ for all $e\gg 0$. Thus for $e$ large, $cdx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}$ implying that $x\in I^{*\mathfrak{a}^{t}}$.
@@
    This proof is nearly identical to the tight closure case, using [SH06, Corollary 6.8.12]. ∎
In [HY03, Proposition 1.3(4)], the authors prove that the Hara-Yoshida $\mathfrak{a}$-tight closure of an ideal $I$ is unaffected by replacing $\mathfrak{a}$ with any ideal with the same integral closure. We prove that the same holds for Vraciu $\mathfrak{a}$-tight closure and $\mathfrak{a}$-integral closure as well.
@@
Let $I,\mathfrak{a},\mathfrak{b}\subset R$ be ideals and let $\mathfrak{b}\subset\mathfrak{a}$ be a reduction. Then ${}^{\mathfrak{a}^{t}}I^{*}=$ ${}^{\mathfrak{b}^{t}}I^{*}$ and $\overline{I}^{\mathfrak{a}^{t}}=\overline{I}^{\mathfrak{b}^{t}}.$
@@
$cdx^{p^{e}}\mathfrak{b}^{\lceil tp^{e}\rceil}\subset cdx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset dI^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil+k}\subset I^{[p^{e}]}\mathfrak{b}^{\lceil tp^{e}\rceil}$
@@
We now turn to the adjusted integral closures. Since $\mathfrak{b}\subset\mathfrak{a},$ from 3.3(3), we know that $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{I}^{\mathfrak{b}^{t}}.$ Let $x\in\overline{I}^{\mathfrak{b}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{b}^{\lceil tn\rceil}\subset I^{n}$ for all $n$ large. Thus for $n$ large, we have
@@
One of the most important theorems involving the ordinary tight and integral closures of ideals is the Briançon-Skoda theorem, see [HH90, Theorem 5.4]. A similar version involving the Frobenius closure is proved in [HH94, Discussion 7.6(c)] or [KZ19, Theorem 2.2]. We note that these adjusted closure operations also satisfy a similar theorem.
Theorem 3.19 (A Briançon-Skoda theorem for the adjusted closures).
Let $\mathfrak{a},I$ be ideals in a ring $R$. Let $I$ be generated by $d$ elements and let $t\in\mathbb{R}_{>0}$.
@@
    If $d=0,$ that is, if $I$ is the zero ideal, then by definition,
    $\overline{(0)^{m}}^{\mathfrak{a}^{t}}=\overline{(0)}^{\mathfrak{a}^{t}}=(0)^{*\mathfrak{a}^{t}}=((0)^{m+1})^{*\mathfrak{a}^{t}}.$
    Now let $I$ be a non-zero ideal and let $x\in\overline{I^{d+m}}^{\mathfrak{a}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n},$ for all $n$ large. Write $n$ as $p^{e}.$ From [Tay21, Lemma 2.1], we know that $I^{(d+m)p^{e}}\subset(I^{[p^{e}]})^{m+1}.$ Thus, we have $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{m+1})^{[p^{e}]}$ for all $e$ large so that $x\in(I^{m+1})^{*\mathfrak{a}^{t}}$.
@@
    Let $\sqrt{(0)}$ denote the nilradical of $R.$ By passing to the quotient ring $R/\sqrt{(0)}$, we may assume that $R$ has no embedded primes. In other words, every $c\in R^{\circ}$ is a nonzerodivisor. Let $x\in\overline{I^{d+m}}^{\mathfrak{a}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n},$ for all $n$ large implying that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n}\cap(c)R$. Using Artin-Rees Lemma, there exists $n^{\prime}$ such that for all $n$ large,
@@
    As $c$ is a nonzerodivisor, we get $x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n-n^{\prime}}$ for all $n$ large. Write $n$ as $p^{e}.$ From [Tay21, Lemma 2.1], we know that $I^{(d+m)p^{e}-n^{\prime}}\subset(I^{[p^{e}]})^{\lceil m-(n^{\prime}/p^{e})\rceil}=(I^{[p^{e}]})^{m},$ for all $e$ large. Thus, $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{m})^{[p^{e}]}$ for all $e$ large implying that $x\in(I^{m})^{F\mathfrak{a}^{t}}.$

### v3

@@ 4. Comparing the new closures
Later in Section 6 (6.10), we see that the inclusions above are generally strict. We note that each of the adjusted closure operations are insensitive to replacing the base ideal $I$ with its corresponding unadjusted ideal closure operation.
@@
Let $I,\mathfrak{a}\subset R$ be ideals of positive height and $t\in\mathbb{R}_{>0}$. Then, we have the following equalities.
@@
    $(I^{F})^{F\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$ and $^{\mathfrak{a}^{t}}(I^{F})^{F}=\,^{\mathfrak{a}^{t}}I^{F}$, and
@@
    $(I^{*})^{*\mathfrak{a}^{t}}=I^{*\mathfrak{a}^{t}}$ and $^{\mathfrak{a}^{t}}(I^{*})^{*}=\,^{\mathfrak{a}^{t}}I^{*}$, and
@@
    $\overline{J}^{\mathfrak{a}^{t}}=\overline{I}^{\mathfrak{a}^{t}}$, where $J=\overline{I}$. In particular, if $L\subset I$ is a reduction, then $\overline{L}^{\mathfrak{a}^{t}}=\overline{I}^{\mathfrak{a}^{t}}.$
@@
Since $^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}={}^{\mathfrak{a}^{t}}I^{F}$ and $^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{*})^{*}={}^{\mathfrak{a}^{t}}I^{*},$ it is easy to check the equalities using the inclusions $I\subset I^{F}\subset{}^{\mathfrak{a}^{t}}I^{F}$ and $I\subset I^{*}\subset{}^{\mathfrak{a}^{t}}I^{*},$ so we turn to the Hara-Yoshida versions.
@@
    We will show $(I^{F})^{F\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$. As $I\subset I^{F}$, we have $I^{F\mathfrak{a}^{t}}\subset(I^{F})^{F\mathfrak{a}^{t}}$. To see the reverse containment, let $x\in(I^{F})^{F\mathfrak{a}^{t}}$. Then $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{F})^{\left[p^{e}\right]}=I^{\left[p^{e}\right]}$ for all $e$ large. Thus, $x\in I^{F\mathfrak{a}^{t}}$.
@@
    We will show $(I^{*})^{*\mathfrak{a}^{t}}=I^{*\mathfrak{a}^{t}}$. As $I\subset I^{*}$, we have $I^{*\mathfrak{a}^{t}}\subset(I^{*})^{*\mathfrak{a}^{t}}$. To see the reverse containment, let $x\in(I^{*})^{*\mathfrak{a}^{t}}$. Then there exists $c\in R^{\circ}$ such that $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{*})^{\left[p^{e}\right]}$ for all $e$ large.
    For the moment, write $I^{*}=(x_{1},\ldots,x_{n})$. For each $i$ there is a $c_{i}\in R^{\circ}$ such that $c_{i}x_{i}^{p^{e}}\in I^{\left[p^{e}\right]}$ for all $e\gg 0$, and so taking $d=\prod_{i}c_{i}$, for all $e\gg 0$ and $y\in I^{*}$, $dy^{p^{e}}\in I^{\left[p^{e}\right]}$. Hence $d(I^{*})^{\left[p^{e}\right]}\subset I^{\left[p^{e}\right]}$ for all $e\gg 0$. Thus for $e$ large, $cdx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}$ implying that $x\in I^{*\mathfrak{a}^{t}}$.
@@
    This proof is nearly identical to the tight closure case, but we include it for completeness. First, as $I\subset J$, we have $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{J}^{\mathfrak{a}^{t}}$ for any $t$. To see the reverse containment, let $x\in\overline{J}^{\mathfrak{a}^{t}}$. Then, there is a $c\in R^{\circ}$ such that for all $n\gg 0$, $cx^{n}\mathfrak{a}^{\lceil nt\rceil}\subset J^{n}$.
    For the moment, write $J=(x_{1},\ldots,x_{m})$. As $J=\overline{I},$ by [SH06, Corollary 6.8.12] and Proposition 3.2, we have for each $i$ a $c_{i}\in R^{\circ}$ such that $c_{i}x_{i}^{n}\in I^{n}$ for all $n\geq 0$. So, by taking $d=\prod_{i}c_{i}$, we have for any $y\in J^{n}$, $dy\in I^{n}$ for all $n\geq 0$, namely $dJ^{n}\subset I^{n}$ for all $n\geq 0$. Hence, $dcx^{n}\mathfrak{a}^{\lceil nt\rceil}\subset dJ^{n}\subset I^{n}$ for all $n\gg 0$, so $x\in\overline{I}^{\mathfrak{a}^{t}}$. ∎
In [HY03, Proposition 1.3(4)], the authors prove that the Hara-Yoshida $\mathfrak{a}$-tight closure of an ideal $I$ is unaffected by replacing $\mathfrak{a}$ with any ideal with the same integral closure. We prove that the same holds for Vraciu $\mathfrak{a}^{t}$-tight closure and $\mathfrak{a}^{t}$-integral closure as well.
@@
Let $R$ be a ring, $I,\mathfrak{a},\mathfrak{b}\subset R$ be ideals, and suppose $\mathfrak{b}\subset\mathfrak{a}$ is a reduction. Then $^{\mathfrak{a}^{t}}I^{*}=$ $^{\mathfrak{b}^{t}}I^{*}$ and $\overline{I}^{\mathfrak{a}^{t}}=\overline{I}^{\mathfrak{b}^{t}}.$
@@
$cdx^{p^{e}}\mathfrak{b}^{\lceil tp^{e}\rceil}\subset cdx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset dI^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\mathfrak{a}^{\lceil tp^{e}\rceil+k}\subset I^{[p^{e}]}\mathfrak{b}^{\lceil tp^{e}\rceil}$
@@
We now turn to the $\mathfrak{a}^{t}$-integral closure. Since $\mathfrak{b}\subset\mathfrak{a},$ from 3.3(3), we know that $\overline{I}^{\mathfrak{a}^{t}}\subset\overline{I}^{\mathfrak{b}^{t}}.$ Let $x\in\overline{I}^{\mathfrak{b}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{b}^{\lceil tn\rceil}\subset I^{n}$ for all $n$ large. Thus for $n$ large, we have
@@
As for the ordinary tight and integral closures, the adjusted versions agree for principal ideals.
Proposition 4.4.
Let $R$ be a ring, $\mathfrak{a}\subset R$ an ideal of positive height, and $t\in\mathbb{R}_{>0}$. Then, for any $x\in R^{\circ}$, $(x)^{*\mathfrak{a}^{t}}=\overline{(x)}^{\mathfrak{a}^{t}}$ and $^{\mathfrak{a}^{t}}(x)^{*}=\overline{(x)}$.
Proof.
Since the $\mathfrak{a}^{t}$-tight closure is always contained in the $\mathfrak{a}^{t}$-integral closure, it suffices to prove the reverse containment. Let $y\in\overline{(x)}^{\mathfrak{a}^{t}}$, so that there exists $c\in R^{\circ}$ such that for all $n\gg 0,$ $cy^{n}\mathfrak{a}^{\lceil tn\rceil}\subset(x)^{n}$. Therefore, for $e\gg 0$, we may take $n=p^{e}$ and so
$cy^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(x)^{p^{e}}=(x)^{[p^{e}]}$
Therefore $y\in(x)^{*\mathfrak{a}^{t}}$, and so $\overline{(x)}^{\mathfrak{a}^{t}}\subset(x)^{*\mathfrak{a}^{t}}$. The Vraciu version holds similarly, after recalling that the Vraciu $\mathfrak{a}^{t}$-integral closure is the usual integral closure in this case. ∎
4.1. Briançon-Skoda theorems
First shown by Briançon-Skoda in [SBc74] for the ring of convergent power series in $n$ variables over the complex numbers, and then Lipman-Tessier in [LT81] for pseduo-rational local rings, the Briançon-Skoda theorem relates $\overline{I^{n+k}}$ and $I^{n}$. Later, Hochster-Huneke developed a tight closure version ([HH90, Theorem 5.4]) which relates $\overline{I^{n+k}}$ and $(I^{n})^{*}$ using a much simpler proof. A similar version involving the Frobenius closure is proved in [HH94, Discussion 7.6(c)] or [KZ19, Theorem 2.2]. Efforts have been made to uniformize the constant $k$, most notably Huneke in [Hun92] who showed that, under some mild hypotheses, one can take $k$ to be uniform over all ideals $I$. We note that the adjusted closure operations satisfy similar theorems.
Theorem 4.5 (A Briançon-Skoda theorem for the adjusted closure operation).
Let $R$ be a ring, and $\mathfrak{a},I\subset R$ be ideals of positive height. Let $I$ be generated by $d$ elements and let $t\in\mathbb{R}_{>0}$.
@@ 4.1. Briançon-Skoda theorems
Further, suppose $(R,\mathfrak{m})$ is local of dimension $\dim R=n$ and $R/\mathfrak{m}$ is infinite.
  1. (c)
    Then, $\overline{I^{n+m}}^{\mathfrak{a}^{t}}\subset(I^{m+1})^{*\mathfrak{a}^{t}}$ for all $m\geq 0$, and
  2. (d)
    $\overline{I^{n+m}}^{\mathfrak{a}^{t}}\subset(I^{m})^{F\mathfrak{a}^{t}}$ for all $m\geq 1$.
@@
    Let $x\in\overline{I^{d+m}}^{\mathfrak{a}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n},$ for all $n$ large. In particular, for all $n=p^{e}$ with $e\gg 0$, from [Tay21, Lemma 2.1], we know that $I^{(d+m)p^{e}}\subset(I^{[p^{e}]})^{m+1}.$ Thus, we have $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{m+1})^{[p^{e}]}$ for all $e$ large so that $x\in(I^{m+1})^{*\mathfrak{a}^{t}}$.
@@
    We first assume that $R$ is reduced and thus has no embedded primes. In other words, every $c\in R^{\circ}$ is a nonzerodivisor. Let $x\in\overline{I^{d+m}}^{\mathfrak{a}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n},$ for all $n$ large implying that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n}\cap(c)R$. Using the Artin-Rees lemma, there exists $n^{\prime}$ such that for all $n$ large,
@@
    As $c$ is a nonzerodivisor, we get $x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{(d+m)n-n^{\prime}}$ for all $n$ large. As before, taking $n=p^{e}$, from [Tay21, Lemma 2.1], we know that $I^{(d+m)p^{e}-n^{\prime}}\subset(I^{[p^{e}]})^{\lceil m-(n^{\prime}/p^{e})\rceil}=(I^{[p^{e}]})^{m},$ for all $e$ large. Thus, $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(I^{m})^{[p^{e}]}$ for all $e$ large implying that $x\in(I^{m})^{F\mathfrak{a}^{t}}.$
@@
  3. (c)
    Let $\ell$ be the analytic spread of $I$, and we know $\ell\leq n$. Then, under our hypotheses we have that $I$ has a minimal reduction $J$ generated by $\ell$ elements by [SH06, Prop. 8.3.7], say $J=(x_{1},\ldots,x_{\ell}).$ Then $\overline{J^{m}}=\overline{I^{m}}$ for any $m\geq 0$. Hence, $\overline{I^{n+m}}^{\mathfrak{a}^{t}}=\overline{J^{n+m}}^{\mathfrak{a}^{t}}$ by Proposition 4.2(c). By part (a) of this theorem, we have
    $\overline{J^{m+n}}^{\mathfrak{a}^{t}}\subset(J^{m+(n-\ell)+1})^{*\mathfrak{a}^{t}}\subset(J^{m+1})^{*\mathfrak{a}^{t}}\subset(I^{m+1})^{*\mathfrak{a}^{t}}$
    since $J^{m+1}\subset I^{m+1}$, which shows the result. The proof of (d) is nearly identical. ∎